OpenStudy (anonymous):

Using Hamilton’s definition of complex numbers, show that complex addition is both associative and commutative. Show that complex multiplication is both associative and commutative.

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

@Hunus

terenzreignz (terenzreignz):

What in the world is Hamilton's definition?.... Stand by... googling...

terenzreignz (terenzreignz):

Can't find it... you better tell me what Hamilton's definition is... All I can find are quaternions...

terenzreignz (terenzreignz):

Or... better call the big guns in :D @satellite73 ?

OpenStudy (anonymous):

@terenzreignz ^here are my notes on Hamilton from my History of Math course. Quaternions are mentioned in the notes also.

terenzreignz (terenzreignz):

okay... it seems, and correct me if I'm wrong (and if I'm wrong you should really be able to spot it... assuming you read your notes... I just skimmed through it) The definition for complex numbers is \[\huge \mathbb{C} = \left\{a+bi \left| \quad a,b\in \mathbb{R} \quad , \quad i = \sqrt{-1}\right.\right\}\]

OpenStudy (anonymous):

haha well yeah I know that in general from understanding complex numbers @terenzreignz

terenzreignz (terenzreignz):

No need to keep tagging me... just call me TJ, and we'll get along ;) So, let's so associativity, and let \[\LARGE \color{green}{z_1 , z_2, z_3} \in \mathbb{C}\] \[\Large z_1 =a_1+b_1i\]\[\Large z_2 = a_2+b_2i\]\[\Large z_3 = a_3+b_3i\]

terenzreignz (terenzreignz):

*show* associativity

terenzreignz (terenzreignz):

\[\Large (z_1+z_2)+z_3=[(a_1+b_1i)+(a_2+b^2i)]+a_3+b_3i\]

terenzreignz (terenzreignz):

\[\Large = [(a_1+a_2)+(b_1+b_2)i]+a_3+b_3i\]\[\Large =(a_1+a_2+a_3)+(b_1+b_2+b_3)i\] \[\Large = a_1+b_1i + [(a_2+a_3)+(b_2+b_3)i]\]\[\Large = a_1+b_1i+[(a_2+b_2i)+(a_3+b_3i)]\] \[\Large \color{blue}{=z_1+(z_2+z_3)}\] Thus showing associativity

terenzreignz (terenzreignz):

Maybe the commutativity bit, you can do yourself? :)

OpenStudy (anonymous):

Okay. Thank you. So for commutativity, do I just show some combination of z1 + z2 + z3 equaling another combination of z1 + z2 + z3?

terenzreignz (terenzreignz):

Just z1 and z2 should be enough. Just need to show that z1 + z2 = z2+ z1 and that's enough :)

OpenStudy (anonymous):

Okay. Cool. And for the multiplication .. should they reflect the addition associativity and commutativity?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Yes, they should :)

OpenStudy (anonymous):

Okay. Thank you